40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 14398     Entry time: Mon Jan 14 10:06:53 2019
Author: gautam 
Type: Update 
Category: VAC 
Subject: Vent 82 complete 

[chub, gautam]

  • IFO pressure was ~2e-4 torr when we started, on account of the interlock code closing all valves because the N2 line pressure dropped below threshold (<65 psi)
  • Chub fixed the problem on the regulator in the drill-press area where the N2 tanks are, the N2 line is now at ~75 psi so that we have the ability to actuate valves if we so desire
  • We decided that there is no need to vent the pumpspool this time - avoiding an unnecessary turbo landing, so the pumpspool is completely valved off from the main volume and TPs 1-3 are left running
  • Went through the pre-vent checklist:
    • Chub measured particle count, deemed it to be okay (I think we should re-locate the particle counter to near 1X8 because that is where the air enters the IFO anyways, and that way, we can hook it up to the serial device server and have a computerized record of this number as we had in the past, instead of writing it down in a notebook)
    • Checked that the PSL was manually blocked from entering the IFO
    • Walked through the lab, visually inspected Jam Nuts and window covers, all was deemed okay
  • Moved 2 tanks of N2 into the lab on account of the rain
  • Started the vent at ~930am PST
    • There were a couple of short bursty increases in the pressure as we figured out the right valve settings but on average, things are rising at approx the same rate as we had in vent 81...
    • There was a rattling noise coming from the drypump that is the forepump for TP2 (Agilent) - turned out to be the plastic shell/casing on the drypump, moreover, the TP2 diagnostics (temperature, current etc) are all normal.
    • The CC1 gauge (Hornet) is supposed to have an auto-shutoff of its High Voltage when the pressure exceeds 10 mTorr, but it was reporting pressures in the 1 mTorr range even when the adjacent Pirani was at 25 torr. To avoid risk of damage, we manually turned the HV off. There needs to be a python script that can be executed to transition control between the remote and local control modes for the hornet, we had to Power Cycle the gauge because it wouldn't give us local control over the HV.
    • Transitioned from N2 to dry air at P1a=25 torr. We had some trouble finding the correct regulator (left-handed thread) for the dry air cylinders, it was stored in a cabinet labelled green optics no
    • Disconnected dry air from VV1 intake once P1b reached 700 torr, to let lab air flow into the IFO and avoid overpressuring.
    • VA* and VAV* valves were opened so as to vent the annuli as we anticipate multiple chamber openings for this vent.

As of 8pm local time, the IFO seems to have equilibriated to atmospheric pressure (I don't hear the hiss of in-rushing air near 1X8 and P1a reports 760 torr). The pumpspool looks healthy and there are no signs in the TP diagnostics channels that anything bad happened to the pumps. Chub is working on getting the N2 setup more robust, we plan to take the EY door off at 9am tomorrow morning with Bob's help.

* I took this opportunity to follow instructions on pg 29 of the manual and set the calibration for the SuperBee pirani gauge to 760 torr so that it is in better agreement with our existing P1a Pirani gauge. The correction was ~8% (820-->760).

Attachment 1: Vent82Summary.png  1.283 MB  Uploaded Mon Jan 14 20:08:37 2019  | Hide | Hide all
Vent82Summary.png
ELOG V3.1.3-